cos^2(θ)tan^2(θ)
1
Expert's answer
2020-11-17T15:42:55-0500

Recall that "\tan\theta=\frac{\sin\theta}{\cos\theta}"

"\cos^2\theta \tan^2\theta=\cos^2\theta.\left(\frac{\sin\theta}{\cos\theta}\right)^2\\
=\cos^2\theta.\frac{\sin^2\theta}{\cos^2\theta}=\sin^2\theta."


Therefore, "\cos^2\theta \tan^2\theta=\sin^2\theta."